# Closed-loop Long-horizon Robotic Planning via Equilibrium Sequence Modeling

### v4

@@ B.3 Implementation details
Prompt. Our approach involves two LLMs, one LLM as the equilibrium planner and an additional LLM as an optional world model. The planner’s prompt is illustrated in Fig. 9. As can be seen, it consists of system prompt, task definition, environment description, history feedback, and draft plan. Notably, the system prompt is modified from Hu et al. (2024), and the environment section describes the initial environment sorted by rooms, including the object names with their IDs within each room. For the additional world model, we adopt a similar prompt, except that it receives more information about the initial environment, including edges in the scene graph that correspond to spatial relations The world model is then asked to predict the environmental feedback given a generated plan.
@@
Finetuning. Both our equilibrium planner and the world model are finetuned in a supervised manner. The equilibrium planner is finetuned for 6 iterations with a learning rate of 0.0002. At each iteration, we curate training data by pairing fixed points and ground truths, where the fixed points are sampled from an equilibrium memory of past iterations. To prevent overfitting to the history equilibrium, a decay ratio of 0.5 is used when sampling from the fixed points of previous iterations. Thereafter, we update the model parameters using gradient descent according to Eq. 6 for one epoch per iteration. For the world model, we collect all interacting experiences between the planner and the environment, including plans and feedback, and finetune it for 5 epochs using the same learning rate of 0.0002.
@@
Inference. The inference procedure of our planner was described in Algorithm 1, which involves a nested equilibrium solving process. Given the environment and the task instruction, we initialize the draft plan $x_{0}$ and the feedback $c_{0}$ as null and iterate through a nested loop. Each inner loop reuses the feedback from the outer loop to self-refine the draft plan, and after it converges, we update the feedback by interacting with the environment or world model. This process continues until it reaches an upper bound on the outer loop, which we set to 10 to match Tree-Planner (Hu et al., 2024). We adopt greedy sampling for the LLM, except that the first refinement step uses top-k sampling with $k=10$ for higher diversity. Since most of the prompt remains unchanged during equilibrium solving, we implement KV cache to accelerate the inference process, which can be further improved with parallel decoding techniques (Santilli et al., 2023; Cai et al., 2024; Kou et al., 2024).
@@
Action fail, trackback
@@
Task fail, trackback
@@
Task fail, trackback
@@
Figure 11: Visualization of our self-correction process with different types of feedback. The external and internal feedback refer to the feedback from the environment and the world model, respectively. The task instruction is “Open email application, open new emails and respond accordingly”.
@@ Appendix C Additional Results
To further illustrate the effectiveness of equilibrium sequence modeling in closed-loop planning, we additionally exemplify our self-correction process on a long-horizon robotic planning task in Fig. 10. Compared to SELF-REFINE (Madaan et al., 2023) and Tree-Planner (Hu et al., 2024), our approach is more competent in revising a long plan through few forward passes without additional system 2. This is attributed to our efficient training scheme for teaching planners to self-refine.
We also compare different types of feedback utilized by our planner in Fig. 11. As can be observed, internal feedback alone cannot enable successful replanning, but it can reduce environmental interactions prior to convergence. This confirms the effectiveness of both internal and external feedback in closed-loop planning and thus justifies the framework design of our equilibrium planner.
@@ Appendix D Limitations
While our equilibrium sequence modeling improves the planning capability of LLMs, we identify the following failure scenarios during the experiments: (1) hallucination of the equilibrium planner and the world model as in vanilla LLMs; (2) lack of awareness of history context such as previously grabbed objects. The latter may be resolved with reasoning techniques as in Yao et al. (2023b).
In a broader sense, our method may be limited in generalizing to new domains because it requires the ground truth and environmental feedback during training. These procedures with the equilibrium solving process results in lower training efficiency. Also, the current formulation only considers the explicit output plan without implicit reasoning steps. Furthermore, our model has only text input and no visual input, which limits its applicability in the real world. This can be resolved by introducing video-based planners (Du et al., 2024) and world models (Yang et al., 2024; Brooks et al., 2024).

### v5

@@ B.3 Implementation Details
Prompt. Our approach involves two LLMs, one LLM as the equilibrium planner and an additional LLM as an optional world model. The planner’s input is illustrated in Figure 9. As can be seen, it consists of five parts: system prompt, task definition, environment description, history feedback, and draft plan. Notably, the system prompt is modified from Hu et al. (2024), and the environment section describes the initial environment sorted by rooms, including the object names with their IDs within each room. For the optional world model, we adopt a similar prompt, except that it receives more information about the initial environment, including edges in the scene graph that indicate spatial relations. This additional information helps the world model to better predict the environmental feedback given a generated plan.
@@
Finetuning. Both our equilibrium planner and the world model are finetuned in a supervised manner. The equilibrium planner is finetuned for 6 iterations with a learning rate of 0.0002. The training data is constructed adaptively using all previous equilibrium solutions. Specifically, an equilibrium memory is maintained that buffers all equilibrium solutions, including the newest ones. At each iteration, we curate the training data by weighted sampling from this memory (where the newest solutions are sampled more frequently) and then pairing them with the ground truths. To prevent overfitting to the history equilibrium, a decay ratio of 0.5 is used when sampling from the fixed points of previous iterations. Thereafter, we update the model parameters using gradient descent according to Equation 6 for one epoch per iteration.
For the world model, we collect all interacting experiences between the planner and the environment, including plans and feedback, and finetune it for 5 epochs using the same learning rate of 0.0002. The world model is initialized from Llama 3 8B Instruct and supervised finetuned on all the planner’s environmental interactions. This procedure takes place after the equilibrium planner has completed training, so that all of its data can be leveraged at once. The finetuning data is constructed in a format similar to Figure 9 in the appendix, with a different order to predict feedback from the plan. Finetuning the world model takes about 30 hours due to its longer context (e.g. spatial relations).
@@
Inference. The inference of our planner is described in Algorithm 1, which involves a nested equilibrium solving process. Given the environment and the task instruction, we initialize the draft plan $x_{0}$ and the feedback $c_{0}$ as null and iterate through a nested loop. Each inner loop reuses the feedback from the outer loop to self-refine the draft plan, and after the inner loop converges, we update the feedback by interacting with the environment or world model. The ratio of environmental interactions to world model calls is currently set to 1:1, i.e., the planner alternates between using the environmental feedback and the world model at each loop. Note that it is possible to reduce this ratio and the number of environmental interactions required if the planner has access to a more accurate world model. This process continues until it converges to an equilibrium point or reaches an upper bound on the outer loop, which we set to 10 to match Tree-Planner (Hu et al., 2024) but is rarely reached. Thus, the inference compute used is mostly determined by the convergence speed of the model itself.
Our LLM sampling strategy facilitates model convergence. Specifically, we use greedy sampling to stabilize LLM outputs, except that the first refinement step uses top-k sampling with $k=10$ for higher diversity. Since most of the text prompt remains unchanged during equilibrium solving, we employ KV cache to accelerate inference, which can be further improved with parallel decoding techniques (Santilli et al., 2023; Cai et al., 2024; Kou et al., 2024).
@@
Action fail, trackback * 2
@@
Task fail, trackback * 6
@@
Task fail, trackback * 2
@@
Output log
[WALK] <home_office>
[WALK] <laptop>
[FIND] <laptop>
[OPEN] <laptop>
[FIND] <mail>
[GRAB] <mail>
[FIND] <desk>
[OPEN] <desk>
[PUTBACK] <laptop> <desk>
[PUTBACK] <mail> <desk>
[FIND] <keyboard>
[TYPE] <keyboard>
Wrong state:
<character> {SITTING}
Wrong relative pos:
<character> and <mail>
<character> and <chair
Same plan, task fail
Output log
[WALK] <home_office>
[WALK] <laptop>
[FIND] <laptop>
[OPEN] <laptop>
Action fail, trackback * 1
[GRAB] <laptop>
[FIND] <chair>
[SIT] <chair>
[SWITCHON] <laptop>
[FIND] <mail>
[GRAB] <mail>
[READ] <mail>
[FIND] <keyboard>
[TYPE] <keyboard>
Task fail, trackback * 3
[FIND] <desk>
[OPEN] <desk>
Action fail, trackback * 7
[WALK] <table>
[OPEN] <laptop>
Action fail, trackback * 2
[WALK] <desk>
[OPEN] <laptop>
Action fail, trackback * 3
[TURNTO] <laptop>
[LOOKAT] <laptop>
[SWITCHON] <laptop>
[FIND] <chair>
[SIT] <chair>
[FIND] <mail>
[GRAB] <mail>
[READ] <mail>
[FIND] <keyboard>
[TYPE] <keyboard>
Task success
Output log
[WALK] <home_office>
[WALK] <laptop>
[FIND] <laptop>
[SWITCHON] <laptop>
[TURNTO] <laptop>
[LOOKAT] <laptop>
[FIND] <mail>
[READ] <mail>
[TURNTO] <mail>
[POINTAT] <mail>
[GRAB] <mail>
[FIND] <keyboard>
[TYPE] <keyboard>
Wrong state:
<character> {SITTING}
Wrong relative pos:
<character> and <mail>
<character> and <chair>
[WALK] <home_office>
[WALK] <laptop>
[FIND] <laptop>
[SWITCHON] <laptop>
[TURNTO] <laptop>
[LOOKAT] <laptop>
[FIND] <chair>
[SIT] <chair>
[FIND] <mail>
[GRAB] <mail>
[READ] <mail>
[TURNTO] <mail>
[POINTAT] <mail>
[WRITE] <mail>
[SWITCHOFF] <laptop>
Invalid command:
[WRITE] <mail>
[WALK] <home_office>
[WALK] <laptop>
[FIND] <laptop>
[SWITCHON] <laptop>
[TURNTO] <laptop>
[LOOKAT] <laptop>
[FIND] <chair>
[SIT] <chair>
[TURNTO] <laptop>
[POINTAT] <laptop>
[TURNTO] <laptop>
[LOOKAT] <laptop>
[FIND] <mail>
[GRAB] <mail>
[READ] <mail>
[TYPE] <laptop>
Task success
(a) SELF-REFINE
(b) Tree-Planner
(c) Ours
Figure 11: Visualization of our self-correction process in comparison with baselines. This example uses only environmental feedback. The task instruction is “Open email application, open new emails and respond accordingly”. Our proposed method succeeds with fewer external feedback.
@@
Figure 12: Visualization of our self-correction process with different types of feedback. The task instruction is “Open email application, open new emails and respond accordingly”.
Output log
[WALK] <bathroom>
[WALK] <cleaning_solution>
[FIND] <cleaning_solution>
[GRAB] <cleaning_solution>
[FIND] <rag>
[GRAB] <rag>
[POUR] <cleaning_solution>
<rag>
[PUTOBJBACK]
<cleaning_solution>
[FIND] <sink>
[TURNTO] <sink>
[LOOKAT] <sink>
[TOUCH] <sink>
Inner self-refinement
[WALK] <bathroom>
[WALK] <cleaning_solution>
[FIND] <cleaning_solution>
[GRAB] <cleaning_solution>
[FIND] <rag>
[POUR] <cleaning_solution>
<rag>
[PUTOBJBACK]
<cleaning_solution>
[FIND] <sink>
[TURNTO] <sink>
[LOOKAT] <sink>
[TOUCH] <sink>
Inner loop convergence
Environmental feedback
Wrong relative pos:
<cleaning_solution> and <sink>
Output log
[WALK] <bathroom>
[WALK] <cleaning_solution>
[FIND] <cleaning_solution>
[GRAB] <cleaning_solution>
[FIND] <sink>
[POUR] <cleaning_solution>
<sink>
[TURNTO] <sink>
[LOOKAT] <sink>
[TOUCH] <sink>
Inner self-refinement
[WALK] <bathroom>
[WALK] <cleaning_solution>
[FIND] <cleaning_solution>
[GRAB] <cleaning_solution>
[FIND] <sink>
[POUR] <cleaning_solution>
<sink>
[WAIT]
[FIND] <rag>
[GRAB] <rag>
[WIPE] <rag> <sink>
[PUTOBJBACK] <rag>
[SWITCHON] <faucet>
[RINSE] <sink>
[SWITCHOFF] <faucet>
Inner loop convergence
Environmental feedback
invalid command:
[WAIT]
Output log
[WALK] <bathroom>
[WALK] <cleaning_solution>
[FIND] <cleaning_solution>
[GRAB] <cleaning_solution>
[FIND] <sink>
[POUR] <cleaning_solution>
<sink>
[WIPE] <sink>
[PUTOBJBACK]
<cleaning_solution>
Inner loop convergence
Task success
[…]
Figure 13: Visualization of our self-correction process with inner-loop traces. This example uses only environmental feedback. The task instruction is “Spray some cleaner and wait a minute then wipe it out”. As can be seen, each inner loop improves planning quality: the first removes a redundant [GRAB] action, while the second adds a crucial [WIPE] action and other interesting details.
Output log
[WALK] <bathroom>
[WALK] <electric_shaver>
[FIND] <electric_shaver>
[GRAB] <electric_shaver>
[FIND] <mirror>
[TURNTO] <mirror>
[LOOKAT] <mirror>
[FIND] <chair>
[SIT] <chair>
[SWITCHON] <electric_shaver>
[FIND] <beard>
[TURNTO] <beard>
[POINTAT] <beard>
[FIND] <beard>
[WIPE] <beard>
Inner loop convergence
invalid command:
[FIND] <beard>
Task fail
Output log
[WALK] <bathroom>
[WALK] <electric_shaver>
[FIND] <electric_shaver>
[GRAB] <electric_shaver>
[FIND] <mirror>
[TURNTO] <mirror>
[LOOKAT] <mirror>
[SWITCHON] <electric_shaver>
[RUN] <mirror>
[TURNTO] <mirror>
[LOOKAT] <mirror>
[SWITCHOFF] <electric_shaver>
[PUTOBJBACK] <electric_shaver>
Inner self-refinement
[WALK] <bathroom>
[WALK] <electric_shaver>
[FIND] <electric_shaver>
[GRAB] <electric_shaver>
[FIND] <mirror>
[TURNTO] <mirror>
[LOOKAT] <mirror>
[SWITCHON] <electric_shaver>
[SWITCHOFF] <electric_shaver>
Inner loop convergence
Task success
(a) Prompt-based self-refinement
(b) Ours
Figure 14: Visualization of our self-correction process in comparison with prompting-based method without any feedback. The task instruction is “Pick up razor and shave yourself”. As can be seen, the prompting-based model converges without any self-correction, while our model achieves success.
@@ Appendix C Additional Results
Effectiveness of our full method. We exemplify the self-correction trajectories of our full method in Figs. 10 and 11. Compared to SELF-REFINE (Madaan et al., 2023) and Tree-Planner (Hu et al., 2024), our approach is more competent in revising a long plan through few forward passes without additional system 2. This is attributed to our efficient training scheme for teaching planners to self-refine. We also compare different types of feedback utilized by our planner in Figure 12. As can be observed, internal feedback alone cannot enable successful replanning, but it can reduce environmental interactions prior to convergence. This confirms the effectiveness of both internal and external feedback in closed-loop planning.
Effectiveness of the inner loop. Table 6 illustrates that even without feedback, our method yields significant performance improvements. To better understand this, we visualize the self-refining traces of the inner loop in Figures 13 and 14. Specifically, Figure 13 shows the inference trace at each inner-loop iteration. Even though the planner only succeeded in later steps, there are consistent quality improvements in its output during the inner-loop introspection without any feedback. Figure 14 compares our inference trace to a prompting-based alternative, where our method shows to be more effective at steering the output toward a correct plan via the inner loop. The working mechanism of our inner loops was mentioned in the introduction, i.e. allowing for bidirectional dependency as well as scaling of inference-time compute.
Comparison with Tree-Planner. As shown in Table 7, our performance is inferior to Tree-Planner in the no-feedback setting because of the difference in refining a single plan with $<$10 iterations instead of generating 25 or 50 candidate plans (giving Tree-Planner a comparative advantage). However, in the other settings that allow feedback, our method outperforms Tree-Planner by incorporating feedback more flexibly. By taking into account feedback through forward passes of LLMs, our method allows arbitrary changes based on the LLMs’ knowledge, and correcting multiple errors in parallel (Figure 10(c)). In contrast, tree-based alternatives require backtracking in a tree, which is costly when correcting an early mistake and does not fully exploit the implicit knowledge in feedback. For example in Figure 11, Tree-Planner ignored the earlier feedback and made the same mistake twice ([OPEN] ⟨laptop⟩).
Table 5: Number of iterations for LLMs to reach a fixed point.
Mean Std Min Max
Original Llama 6.70 2.12 3.22 14.98
Supervised finetuned Llama 2.46 0.88 2.02 3.80
Ours 3.02 1.61 2.32 4.07
Fixed-point convergence. Table 5 shows the number of iterations for an LLM to reach its fixed point, aggregated over 10 initial plan for each of 60 random tasks. All LLMs considered, including the original Llama, the supervised finetuned Llama, and our model, can converge to a fixed point within a few iterations, regardless of the initial sequence. This is partly due to our greedy sampling strategy (described in Section B.3) that reduces the LLMs’ randomness, after which they tend to repeat themselves (please see Figures 4(a), 10(a) and 11(a)) and converge easily.
Table 6: Effectiveness of our method in the no-feedback setting, where it shows clear performance advantages.
Novel Scene and Task Novel Scene Novel Task
Exec. SR GCR Exec. SR GCR Exec. SR GCR
Supervised 93.55 24.19 32.55 96.84 41.05 49.81 95.94 26.07 35.53
SELF-REFINE 72.58 32.26 52.29 74.74 44.21 62.25 65.38 30.98 51.80
Ours 88.71 33.87 59.98 96.79 49.47 66.60 93.80 34.62 59.06
Table 7: Comparison to Tree-Planner with and without world model. Only our method shows a significant improvement.
Novel Scene and Task Novel Scene Novel Task
World model Exec. SR GCR Exec. SR GCR Exec. SR GCR
Tree-PlannerN=25 $\checkmark$ 96.25 38.71 58.71 98.81 51.58 63.94 96.66 38.46 57.40
Tree-PlannerN=50 $\checkmark$ 96.16 37.10 57.53 98.22 54.74 69.64 96.80 39.32 58.84
Ours 88.71 33.87 59.98 96.79 49.47 66.60 93.80 34.62 59.06
Ours $\checkmark$ 90.32 40.32 65.40 95.79 65.26 79.47 93.38 41.88 62.76
@@ Appendix D Limitations
While our equilibrium sequence modeling improves the planning capability of LLMs, we identify the following failure scenarios during the experiments: (1) hallucination of the equilibrium planner and the world model as in vanilla LLMs; (2) lack of awareness of history context such as previously grabbed objects. The latter can be resolved with the context module in Kim et al. (2023a) or reasoning techniques as in Yao et al. (2023b).
In a broader sense, our method may be limited in generalizing to new domains because it requires the ground truth and environmental feedback during training. These procedures with the equilibrium solving process results in lower training efficiency. Also, the current formulation only considers the explicit output plan without implicit reasoning steps. While it is possible to synergize planning and reasoning with their combined advances, we leave this to future work due to feasibility constraints for small-scale experiments with large reasoning LLMs (Jaech et al., 2024; Guo et al., 2025). Furthermore, our model has only text input and no visual input, which limits its applicability in the real world. This can be resolved by introducing video-based planners (Du et al., 2024) and world models (Yang et al., 2024; Brooks et al., 2024).